### v2

# Advancing Sequential Numerical Prediction in Autoregressive Models

Abstract
Autoregressive models have become the de facto choice for sequence generation tasks, but standard approaches treat digits as independent tokens and apply cross-entropy loss, overlooking the coherent structure of numerical sequences. This paper introduces Numerical Token Integrity Loss (NTIL) to address this gap. NTIL operates at two levels: (1) token-level, where it extends the Earth Mover’s Distance (EMD) to preserve ordinal relationships between numerical values, and (2) sequence-level, where it penalizes the overall discrepancy between the predicted and actual sequences. This dual approach improves numerical prediction and integrates effectively with LLMs/MLLMs. Extensive experiments show significant performance improvements with NTIL. All resources are available at https://github.com/xfey/NTIL.
Advancing Sequential Numerical Prediction in Autoregressive Models
Xiang Feithanks: Equal Contribution
    †Corresponding Author
1   Jinghui Lu∗1   Qi Sun∗2   Hao Feng†1   Yanjie Wang1
Wei Shi1   An-lan Wang1   Jingqun Tang1   Can Huang†1 1ByteDance Inc.   2City University of Hong Kong
{feixiang.77, lujinghui, fenghao.2019}@bytedance.com {wangyanjie.prince, shiwei.11, wanganlan}@bytedance.com {tangjingqun, can.huang}@bytedance.com qisun.new@gmail.com
1 Introduction
In recent years, sequence generation has become a crucial approach for implementing a broad range of AI applications, including visual question answering (Wang et al., 2024d; Reich and Schultz, 2024; Fan et al., 2024; Liu et al., 2024b; He et al., 2025; Wang et al., 2025; Lu et al., 2024b), key information extraction (Kim et al., 2024; Yu et al., 2024; Kang et al., 2024; Wang et al., 2024a; Yi et al., 2025; Lu et al., 2023a, 2024a), object detection (Wen et al., 2024), math reasoning Zhao et al. (2024), text spotting (Li et al., 2024; Yu et al., 2025), and automatic audio recognition (Zhou et al., 2024).
Figure 1: Sequence-level digit token loss illustration.
Autoregressive models, especially large language models (LLMs) such as GPT Achiam et al. (2023), LLaMA Touvron et al. (2023); Dubey et al. (2024), Qwen Yang et al. (2024); Wang et al. (2024c) series, with multi-modal large language models (MLLMs) based on them, now dominate the sequence generation tasks. During training, these models generate sequences token-by-token, typically using cross-entropy (CE) loss, to minimize the negative log-likelihood of the ground truth token at each time step. However, CE loss has several inherent limitations when predicting numerical values. Specifically, CE suffers from Limitation 1 that it ignores the inherent closeness between numerical tokens, where each digit in a numerical prediction is not independent but related to its neighboring digits. For example, in Figure 2(a) and 2(b), for the ground truth token “3”, the CE loss yields same values of $-log(0.5)$ for different prediction distributions. However, the distribution in Figure 2(b) is more accurate, as it assigns higher probability to the neighboring token “2”.
CE also suffers from Limitation 2 that it fails to capture the holistic numerical error when sequential tokens are involved, as it focuses on the precision of each token rather than the overall value. In an autoregressive generation manner, producing a numerical value typically requires consecutive time steps. For example, the target value “0.98” requires the prediction of four sequential tokens — “0”, “.”, “9”, “8”. Thus, a prediction such as 1.01 (“1”,“.”,“0”,“1”) incurs a high CE loss as the first, third and fourth tokens are significantly different from the target tokens. Conversely, a prediction like 1.98 (“1”,“.”,“9”,“8”) could yield a lower CE loss due to a closer match at the token level, despite the overall numerical difference being larger (1.00 vs. 0.03). This discrepancy shows the limitation of CE in evaluating predictions holistically.
Figure 2: Cross-entropy fails to distinguish predictions, whereas EMD correlates smaller loss for better predicted distributions.
To overcome the above issues, we introduce a novel sequence-level numerical prediction loss: Numerical Token Integrity Loss (NTIL). At the token-level, NTIL replaces the traditional CE loss with Earth Mover’s Distance (EMD) (Rubner et al., 1998). Additionally, we enhance the EMD with an Exponential Position-based Weighting scheme (Section 3.1), which leverages place-value number systems to better capture the nuanced differences between numerical distributions at each time step. At the sequence-level, NTIL evaluates the overall numerical difference between predicted and actual sequences through Multi-token Numerical Optimization (Section 3.2), considering all time steps holistically, as illustrated in Figure 1. It enables NTIL to effectively model the actual value of digit sequences, and capture discrepancies across the consecutive numerical range, moving beyond simple token-by-token comparison.
To the best of the authors’ knowledge, it is the first time that EMD is used as an optimization method for autoregressive models. Moreover, our holistic approach is the first of its kind to improve sequential numerical prediction by considering numerical tokens across multiple time steps. Our method can be seamlessly integrated into both LLMs and MLLMs. Experimental results show that NTIL boosts performance in tasks requiring precise numerical outputs, such as object detection, text spotting, and math reasoning (Section 4).
2 Related Work
The Earth Mover’s Distance (EMD) measures the minimal cost of transforming one distribution into another, and has become a valuable metric in deep learning applications. Notably, Wasserstein GAN Arjovsky et al. (2017) uses EMD as its loss function to stabilize training in GANs. Cuturi (2013) and  Courty et al. (2016) also adopted EMD for smoothing the training procedure. Despite the success of EMD, it has not been applied to autoregressive models. Most recently, autoregressive models, especially LLMs, have advanced NLP (Radford et al., 2019; Touvron et al., 2023; Lu et al., 2023b; Cui et al., 2025), and multi-modal tasks (Alayrac et al., 2022; Wang et al., 2024c; Feng et al., 2025; Lu et al., 2025). While the tasks mentioned above require high precision in numerical value prediction, none of the previous works have specifically optimized for this criterion. Our work addresses this gap by focusing on advancing the sequential numerical prediction for autoregressive models.
3 Method
This section details the components of the proposed method. Section 3.1 proposes exponential weighted EMD to single digits; Section 3.2 describes how we go through multiple digital tokens to derive a simple yet effective numerical measure.
3.1 Exponential Position-Based Weighting
For token-level prediction, to address Limitation 1 in Section 1, we replace the conventional CE loss with EMD to account for the ordinal relationship during optimization. The preliminaries for both CE and EMD objectives, and the simplification via numerical prediction, are outlined in Appendix D.
Furthermore, we extend EMD to account for the place-value number systems, where leading digits have greater numerical significance. We implement an exponential weighting scheme to progressively assign weights based on digit positions, to scale their contributions to the loss accordingly:
$\mathbf{W_{exp}}=\left[(1+\sigma)^{n-i-1}\right]_{i=0}^{n-1},$ (1)
where $\sigma$ is the exponential increment rate, and $n$ is the length of consecutive digits. This implementation helps the model understand the order relationship between consecutive numbers.
3.2 Multi-Token Numerical Optimization
To overcome Limitation 2 outlined in Section 1, we propose the following procedure and losses. Differentiable Numerical Value Construction. In this step, we construct the complete numerical value from consecutive discrete digital tokens. Figure 3 illustrates how we obtain the digit index from the predicted distribution using $\operatorname{argmax}$ to derive the integer representation. To maintain differentiability, we employ the $\operatorname{Gumbel-softmax}$ approximation with reduced temperature and noise parameters to ensure consistent results. The resulting tensor is element-wise multiplied with positional indices, scaled by the appropriate powers of 10, and aggregated to obtain the final value. For further implementation details, see Appendix C.
Figure 3: Constructing a numerical value from tokens.
Relative Deviation Metric. For numerical comparison, while absolute difference provides a straightforward measure equivalent to L1 loss, we propose a normalized metric defined as:
$\mathcal{L}_{relative}=\frac{\left|X-Y\right|}{\operatorname{max}\left(X,Y\right)+\epsilon},$ (2)
where $X$ is the sequence-level numerical prediction (e.g., “234”) and $Y$ is the ground truth, and $\epsilon$ is a small quantity to avoid division by zero. This normalization ensures consistency across different magnitude ranges.
Magnitude Deviation Metric We also apply a normalized metric on the order of magnitude as:
$\mathcal{L}_{magnitude}=\operatorname{log}\left(\frac{\operatorname{max}(X,Y)}{\operatorname{min}(X,Y)}\right).$ (3)
The objective penalizes the difference in the order of magnitude between two values. For example, given the pairs $(1,10)$ and $(1,100)$, which have similar $\mathcal{L}_{relatvie}$ values $0.90$ and $0.99$, but differ in $\mathcal{L}_{magnitude}$ value: $\log\left(\frac{10}{1}\right)\approx 2.30$ for the first pair and $\log\left(\frac{100}{1}\right)\approx 4.61$ for the second. This results in a larger penalty for greater differences in magnitude. The final formulation of NTIL combines the above loss functions, with tunable hyperparameters to weight their individual contributions.
$\mathcal{L}=\mathbf{W_{exp}}\operatorname{EMD}+\alpha\cdot\mathcal{L}_{relative}+\beta\cdot\mathcal{L}_{magnitude}$ (4)
RefCOCO RefCOCO+ RefCOCOg
Model Method Val TestA TestB Val TestA TestB Val Test Avg
CE 0.839 0.865 0.784 0.740 0.797 0.664 0.792 0.797 0.785
EMD 0.841 0.864 0.796 0.749 0.805 0.669 0.789 0.799 0.789
PaliGemma (3b) Beyer et al. (2024) Ours 0.844 0.873 0.791 0.750 0.812 0.678 0.804 0.802 0.795
CE 0.855 0.880 0.813 0.801 0.843 0.741 0.799 0.816 0.818
EMD 0.856 0.879 0.822 0.798 0.845 0.743 0.798 0.816 0.820
LLaVA-1.5 (7b) Liu et al. (2024a) Ours 0.858 0.885 0.815 0.800 0.853 0.747 0.802 0.817 0.822
CE 0.767 0.796 0.734 0.706 0.757 0.651 0.722 0.731 0.733
EMD 0.779 0.805 0.738 0.719 0.762 0.657 0.721 0.737 0.740
Yi-VL (6b) Young et al. (2024) Ours 0.777 0.808 0.741 0.717 0.770 0.665 0.727 0.743 0.744
CE 0.897 0.928 0.850 0.841 0.896 0.776 0.851 0.867 0.863
EMD 0.889 0.931 0.843 0.838 0.889 0.772 0.853 0.858 0.859
Qwen2-VL (2b) Wang et al. (2024c) Ours 0.898 0.932 0.849 0.844 0.891 0.788 0.858 0.863 0.866
CE 0.892 0.929 0.841 0.842 0.902 0.784 0.843 0.848 0.860
EMD 0.886 0.926 0.834 0.843 0.901 0.768 0.836 0.843 0.855
Qwen2-VL (7b) Wang et al. (2024c) Ours 0.889 0.931 0.840 0.844 0.904 0.786 0.848 0.853 0.862
Table 1: Performance comparison (Acc@0.5) of models on image grounding tasks.
Dataset
Model Method CTW1500 ICDAR1500 TD500 Total-Text Avg
CE 0.220 0.129 0.183 0.259 0.193
EMD 0.314 0.124 0.252 0.307 0.241
PaliGemma (3b) Ours 0.369 0.155 0.257 0.318 0.263
CE 0.682 0.370 0.753 0.673 0.586
EMD 0.668 0.398 0.778 0.678 0.594
Yi-VL (6b) Ours 0.680 0.403 0.752 0.678 0.597
CE 0.786 0.538 0.851 0.827 0.720
EMD 0.786 0.535 0.867 0.808 0.718
Qwen2-VL (2b) Ours 0.776 0.577 0.854 0.835 0.732
CE 0.771 0.648 0.889 0.864 0.764
EMD 0.762 0.625 0.874 0.860 0.751
Qwen2-VL (7b) Ours 0.770 0.669 0.869 0.872 0.770
CE 0.735 0.490 0.821 0.786 0.675
EMD 0.724 0.545 0.840 0.776 0.690
LLaVA-1.5 (7b) Ours 0.739 0.547 0.839 0.791 0.698
Table 2: Performance (Acc@0.5) on scene text detection tasks.
Model Accuracy (%)
CE EMD Ours
Baichuan2 (7b)
Yang et al. (2023)
44.3 46.6 46.9
Qwen2.5 (1.5b)
Team (2024)
40.3 40.7 42.4
LLaMA3 (8b)
Dubey et al. (2024)
61.9 61.8 61.9
Yi (6b)
Young et al. (2024)
53.0 54.6 54.4
MiniCPM3 (4b)
Hu et al. (2024)
66.8 68.2 68.6
Table 3: Performance comparison of accuracies on the arithmetic calculation task.
Metric Method
LLaVA-1.5
Qwen2-VL
Qwen2-VL
Yi-VL
CE 95.1 75.0 81.3 76.2
EMD 95.3 78.7 81.7 75.1
Accuracy (%) $\uparrow$ Ours 98.3 80.5 85.3 87.4
CE 8.52 30.84 32.34 56.58
EMD 7.98 30.78 31.98 54.78
Time gap (minute) $\downarrow$ Ours 4.14 27.72 24.66 26.58
Table 4: Performance of the clock time recognition task.
Dataset Method
Qwen2-vl
Qwen2-vl
LLaVA-1.5
Yi-VL
PaliGemma
CE 0.139 0.184 0.146 0.143 0.097
EMD 0.130 0.188 0.148 0.142 0.088
Mathvision Ours 0.145 0.191 0.146 0.153 0.098
CE 0.248 0.315 0.140 0.187 0.143
EMD 0.262 0.303 0.157 0.192 0.149
Mathvista Ours 0.251 0.300 0.170 0.222 0.157
Table 5: Performance of the math reasoning task.
PaliGemma LLaVA-1.5 Qwen2-VL Yi-VL
Exp Rel Mag Mathvision Mathvista Mathvision Mathvista Clock_Time Clock_Time
$\times$ 0.096 0.137 0.151 0.166 0.798 0.834
$\times$ 0.095 0.137 0.145 0.154 0.790 0.856
$\times$ 0.094 0.142 0.160 0.143 0.816 0.876
0.098 0.157 0.146 0.170 0.853 0.874
Table 6: Ablations on NTIL. Exp: Exponential Position-Based Weighting. REL: Relative Deviation Metric. Mag: Magnitude Deviation Metric.
Figure 4: Results for quantitative analysis.
4 Experiments and Results
This section presents a comprehensive empirical evaluation of the proposed NTIL across various LLMs/MLLMs (Section 4.1). CE (Shannon, 1948) and EMD (Rubner et al., 1998) are chosen as baselines due to their widespread adoption. The evaluation encompasses multiple task domains that focus on numerical prediction including Image Grounding, Scene Text Detection, Clock Time Recognition, Mathematical Reasoning and Arithmetic Calculations. Appendix 5 provides details on tasks, datasets, and evaluation metrics. We also conduct systematic ablation studies to evaluate the critical components of our approach. Implementation details are available in Appendix A.
4.1 Main Results
4.2 Results of MLLMs
Image Grounding As shown in Table 1, our method outperforms both CE and EMD across nearly all datasets and VLM backbones, as evidenced by the overall performance improvements.
Scene Text Detection Table 3 shows that our method improves accuracy across multiple datasets, demonstrating its effectiveness in predicting multiple object coordinates.
Clock Time Recognition Table 4 demonstrates that NTIL surpasses CE and EMD significantly in performance across all model architectures.
Mathematical Reasoning As shown in Table 5, our method outperforms CE and EMD across all datasets, with the most significant improvements seen in the Mathvision dataset using the Qwen2-VL (2b) and in Mathvista with the Yi-VL (6b).
4.3 Results of LLMs
Arithmetic Calculation As shown in Table 3, our method improves accuracy across multiple LLMs, though LLaMA3 shows minimal gains, possibly due to its extensive pre-training. Overall, the majority of cases show that for numerical predictions, while EMD performs comparably or marginally better than CE loss, NTIL consistently delivers superior results in most scenarios. This underscores the effectiveness and generalizability of NTIL.
4.4 Ablation Analysis
Table 6 indicates that incorporating all components of NTIL generally leads to better performance, as evidenced by the highest scores in most metrics when all components are enabled. As an exception, the inclusion of Magnitude leads to worse results in Mathvision for LLaVA-1.5, which indicates the fluctuation of applying Magnitude in some cases.
4.5 Quantitative Analysis
As shown in Figure 4(a), NTIL achieves the lowest absolute errors among all models, indicating more consistent performance compared to CE and EMD. Figure 4(b) and 4(c) illustrate that NTIL produces more accurate predictions with distributions more concentrated around the ground truth. Overall, NITL offers more stability and lower variability. Qualitative examples can be seen in Appendix F.
5 Conclusion
We propose NTIL, which improves numerical prediction accuracy in LLMs at both the token and sequence levels. Experiments show improvement across multiple datasets and models, highlighting effectiveness of NTIL.
Limitations
The limitations of the NTIL include the exponential position-based weighting scheme, while effective in many cases, had limited or negative impact in certain configurations, such as with the Mathvision dataset and LLaVA-1.5 model. Future exploration could focus on refining the exponential position-based weighting scheme with adaptive strategies to address its inconsistent impact.
Furthermore, NTIL introduces additional computational overhead compared to CE loss, resulting in reduced training efficiency. This trade-off between performance improvement and computational cost needs to be considered in practical applications.
Another limitation arises from the tokenization strategies of certain models like LLaMA-3, which encode common multi-digit numbers (e.g., "123") as a single token. Such cases require special handling in NTIL’s implementation, adding complexity to the framework.
Appendix A Implementation
The proposed loss function is incorporated into the model’s training objective through linear combination with a weighting coefficient $\lambda=0.3$. The hyperparameters governing the loss computation are maintained at $\alpha=\beta=\sigma=0.2$ throughout all experiments, unless otherwise specified. All tasks are trained with a learning rate of $10^{-5}$ for fine-tuning. Our experiments are conducted based on a widely used open source training repository [^1] https://github.com/hiyouga/LLaMA-Factory.
Appendix B Dataset
This section provides a detailed description of each task along with the corresponding evaluation metrics. The illustrations for each task are presented in Figure 5.
Figure 5: The illustrations for each task.
Image Grounding. Grounding task aims to output the bounding box of the corresponding object given a description. We compare on the referring expression comprehension (REC) task on RefCOCO Lin et al. (2014), RefCOCO+ Yu et al. (2016) and RefCOCOg Mao et al. (2016) datasets. The Average Accuracy at $\operatorname{IoU}\geq 0.5$ (Acc@0.5) is used as the evaluation metric.
Scene Text Detection. The scene text detection task focuses on detecting text in natural images. We selected several commonly used datasets: TD500 Yao et al. (2012), ICDAR2015 Karatzas et al. (2015), CTW1500 Yuliang et al. (2017) and Total-Text Ch’ng and Chan (2017) for scene text detection tasks. We utilize the identical metric employed in the image grounding task.
Clock Time Recognition. The perception of clock aims to recognize the specific time by images of clocks. We compare the performance of accuracy and time gap on a widely-used TIME gpiosenka (2022) dataset. The output are formatted as the label of “2_55”, as shown in Figure 6. We use overall accuracy as an metric, and additionally count the time gap between the prediction and the ground truth for further evaluation. For example, the time gap between prediction “4_35” and ground truth “6_20” is 1.75 hours.
Mathematical Reasoning. Completing the mathematical reasoning tasks requires models to understand the context and the image of the mathematical field. We select the MathVista Lu et al. (2023c) and MathVision Wang et al. (2024b) datasets to evaluate models. We utilize exact matching accuracy to evaluate math reasoning task.
Arithmetic Calculations. Calculation task involves training LLMs to perform numerical operations accurately. In this task, the “arithmetic_mix” subset from the widely-used mathematics dataset Saxton et al. (2019) is used for training and evaluation, which contains 2M training and 10k test items. In this task, exact matching accuracy is applied as the evaluation metric.
Appendix C Gumbel Softmax
The Gumbel softmax, also known as Concrete Distribution, is a continuous differentiable approximation to categorical sampling. It replaces the non-differentiable $\operatorname{argmax}$ operation with a softmax function and Gumbel noise. Given logits $\pi_{i}$, the Gumbel softmax sample $y_{i}$ is computed as:
$y_{i}=\operatorname{softmax}\left((log(\pi_{i})+g_{i})/\tau\right),$
where $g_{i}$ is the Gumbel noise, which is i.i.d. samples drawn from the $\operatorname{Gumbel}(0,1)$ distribution, and $\tau$ is the temperature parameter.
The Gumbel noise term $g_{i}$ introduces stochasticity into the sampling process, enabling exploration of the probability space while maintaining differentiability. Moreover, using Gumbel noise also works like regularization, which helps provide gradient information near the decision boundary, to improve generalization ability. The temperature parameter $\tau$ controls the sharpness of the distribution: as $\tau$ approaches 0, the samples become more discrete and closer to one-hot vectors, while higher temperatures make the distribution more uniform. In our implementation, we use $\tau=0.1$ to ensure that the results are consistent with the original $\operatorname{argmax}$ results.
Gumbel softmax is differentiable as it replaces the discrete $\operatorname{argmax}$ with a continuous softmax function, allowing gradients to flow through the sampling process during backpropagation. Thus, Gumbel softmax is widely used in scenarios requiring discrete latent variables in neural networks, such as in VAEsJang et al. (2016) or reinforcement learningHuijben et al. (2022); Wan et al. (2020).
Appendix D Preliminaries
This section first briefly introduces the autoregressive decoding process based on cross-entropy in Section D.1, and then compares and analyzes Earth Mover’s Distance (EMD) in Section D.2.
D.1 Autoregressive Prediction with Cross Entropy
Autoregressive models operate through sequential decoding, generating tokens one at a time conditioned on previously generated tokens. For each position, the model outputs a probability distribution across the vocabulary, employing the Softmax function to select the most probable token during training.
In the context of language modeling tasks, cross-entropy loss serves as the fundamental training objective for autoregressive models. This loss function quantifies the divergence between the predicted probability distribution and the ground truth distribution:
$\mathcal{L}=-\sum_{i}p_{i}\log\left(q_{i}\right),$ (5)
Åwhere $p_{i}$ represents the one-hot encoded ground truth distribution, and $q_{i}$ denotes the model’s predicted probability.
While cross-entropy loss effectively minimizes distributional differences between predictions and labels during training, it exhibits a fundamental limitation in autoregressive decoding: the function treats each class independently, disregarding the inherent relationships between different classes. This limitation becomes particularly problematic when modeling numerical sequences where ordinal relationships between values carry semantic significanceHou et al. (2016), as shown in Figure 2.
D.2 Earth Mover’s Distance
To introduce a distance term when calculating the above-mentioned distribution differences, one method is Earth Mover’s Distance (EMD), also known as Wasserstein distance. It is an evaluation based on optimal transport theory, measuring the minimal cost of transforming one distribution into the other:
$\operatorname{EMD}(P,Q)=\min_{\gamma\in\Gamma(P,Q)}\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}\cdot d\left(x_{i},y_{j}\right),$ (6)
where $P=\left\{\left(p_{i},x_{i}\right)\right\}$ and $Q=\left\{\left(q_{i},y_{i}\right)\right\}$ are two discrete distributions, with $p_{i}$ and $q_{j}$ are the masses at the points $x_{i}$ and $y_{j}$, respectively. The transport plans, represented as $\Gamma(P,Q)$, are all possible ways to move the mass, and $\gamma_{ij}$ represents the amount of mass that is transported from $p_{i}$ to $q_{j}$. The distance matrix $d\left(x_{i},y_{j}\right)$ indicates the cost of transporting masses between points $x_{i}$ and $y_{j}$. A widely-used distance matrix $d$ is Euclidean distance.
Since the distance between labels is explicitly considered, predicted values closer to the label are associated with smaller distance terms. Thus, the Earth Mover’s Distance effectively incorporates distance-based weighting. As illustrated in Figure 2, when the distribution is more concentrated around the label, the EMD loss becomes smaller, thereby reflecting the differences between distributions.
D.3 Predicting Digits with EMD
This section presents our approach to refining distance metrics for numerical representation at the digit level. In traditional autoregressive models, cross-entropy loss is typically employed to predict the probability distributions of individual tokens. However, this method treats each numerical digit as an independent entity, disregarding the continuous relationships between numbers. For example, when the target digit is $4$, a model prediction of $3$ should ideally be considered closer to accurate than a prediction of $9$, as it represents a smaller numerical deviation. To address this limitation, we propose incorporating a distance metric that captures these intrinsic numerical relationships more accurately.
Computational Complexity. As established in D.2, Earth Mover’s Distance (EMD) provides a robust measure for distributional distances, making it particularly well-suited for numerical prediction tasks. Prior research has applied EMD to align hidden representations within neural networks, often requiring the transport plan ($\gamma_{ij}$ in Equ. (6)) to be approximated or recalculated dynamically during training. However, the computational demands of EMD present practical challenges, especially in large-scale deep learning applications. Solving the underlying optimization problem in Equ. (6) has a computational complexity of $O\left((n\times m)^{3}\right)$, which can be prohibitive. Regularized EMD Cuturi (2013) addresses this by employing the Sinkhorn-Knopp algorithm to iteratively refine the transport plan $\gamma_{ij}$ in Equ. (6), reducing complexity to $O\left(k\times n\times m\right)$, where each iteration involves an $O(n\times m)$ matrix operation.
Numerical Prediction Optimization with EMD. When estimating the transport plan, the algorithm’s complexity is generally quadratic. However, when restricted to one-dimensional numerical distributions, where the prediction and target values are aligned in position ($i=j$), the transport plan can be simplified to an identity matrix. Thus, Earth Mover’s Distance emerges as a highly suitable metric for capturing digit-level numerical distance, formulated as:
$\operatorname{EMD}(P,Q)=\sum_{i}\left|x_{i}-y_{i}\right|\cdot\left|i-\operatorname{argmax}\left(Q\right)\right|,$ (7)
where the distance matrix $d\left(x_{i},y_{j}\right)=\left|i-\operatorname{argmax}\left(Q\right)\right|$ refers to the index distance of each digit to the label. Given that the predicted probability distribution $P$ is obtained through the softmax transformation, and the ground truth label $Q$ is represented as a one-hot vector, the gradient of EMD with respect to component $x_{i}$ can be expressed as:
$\frac{\partial\operatorname{EMD}}{\partial x_{i}}=\left\{\left|k-1\right|,\left|k-2\right|,...,\left|k-n\right|\right\}.$ (8)
where $k=\operatorname{argmax}\left(Q\right)$ denotes the index of the label element in the one-hot vector. This gradient exhibits an inverse relationship with the proximity between the predicted distribution and the ground truth: as the prediction approaches the true label, the magnitude of the gradient diminishes. This characteristic is particularly advantageous for numerical prediction tasks, as it inherently accounts for the ordinal relationships between numerical classes, and addresses the fundamental limitation of the conventional cross-entropy.
[…]
0 0.780 0.783 0.789 0.789 0.771
0.1 0.777 0.790 0.799 0.789 0.778
0.2 0.782 0.793 0.795 0.785 0.784
0.3 0.788 0.783 0.786 0.785 0.781
0.4 0.774 0.782 0.778 0.781 0.784
Table 7: Ablation studies of hyper-parameters $\alpha$ and $\beta$
0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7
$\sigma$ 0.784 0.785 0.789 0.788 0.783 0.780 0.770 0.763
$\lambda$ 0.741 0.777 0.798 0.795 0.788 0.786 0.787 0.777
Table 8: Ablation studies of the hyper-parameters $\sigma$ (exp-weighting) and $\lambda$ (coefficient)
Model Params CE EMD NTIL
Qwen2.5 2b 0.509 0.517 0.523
MiniCPM3 4b 0.672 0.676 0.704
Yi 6b 0.375 0.383 0.400
LLaMA3 8b 0.638 0.643 0.646
Table 9: Results of the GSM8k mathematical reasoning task.
Appendix E Ablations
We have supplemented our work with comprehensive experiments on the clock time recognition task using Qwen2-VL-7B, as shown in Table 7 and Table 8. We have bolded some of the outperforming results. The results demonstrate that NTIL adapts well to different hyperparameters.
We have also supplemented NTIL with the GSM8k math reasoning dataset. NTIL brings consistent performance improvement on mixed text-numeric tasks, which shows the effectiveness of our method.
Appendix F Qualitative Examples
Visualizations of the outputs of different losses are shown in Figure 6, and the examples are taken from experimental results using LLaVA-1.5. For image grounding task (Figure 6(a)), the task was to predict the location of “horse back left” in an image. The CE loss (blue box) performed poorly, with predictions far from the ground truth. EMD (red box) showed an improvement, capturing spatial features better, while NTIL (green box) provided the most accurate predictions, closely matching the ground truth (black box). Overall, NTIL outperformed both CE and EMD, demonstrating its effectiveness in this task.
Figure 6(b) presents a qualitative comparision for clock time recognition task. In this case, NTIL provides the most accurate prediction of the clock time, correctly identifying 2:55, which matches the ground truth. EMD performs better than CE, predicting 2:50, but it is still slightly off. CE, however, predicts 5:10, a significant deviation. Overall, NTIL outperforms both EMD and CE in predicting the clock time accurately.
(a) Example in Image Grounding. Blue box is CE prediction, red box is EMD prediction, green box is NTIL prediction. Black box is ground truth.
(b) Example in clock time recognition.
Figure 6: Comparisons between CE, EMD and NTIL.